OpenStudy (anonymous):

A cell phone call costs $0.07 per minute for the first 25 minutes and $0.10 per minute for each additional minute.how much would a 47 minute call cost?

OpenStudy (anonymous):

so the first 25mins is charged at $0.07 per minute so the remaining 22 minutes will be charged at $0.10 so the total cost of 47 min = $0.07(25)+$0.10(22) = $1.75+$2.20 = $3.95

OpenStudy (anonymous):

\[25*\frac{7}{100}+22*\frac{10}{100}=\frac{79}{20}=3.95 \]